# Long-Tailed Continual Learning For Visual Food Recognition

### v1

@@ VI Conclusion
In this work, we focus on visual food recognition within the context of long-tailed continual learning, which is closely related to various real world food based applications. We first introduce two benchmark long-tailed food image datasets including VFN-INSULIN and VFN-T2D, which are generated based on VFN to exhibit the real life food consumption frequency for groups with diabetes. In addition, we propose a novel end-to-end framework that is capable of learning new food classes incrementally in long-tailed data distribution without forgetting the learned knowledge. The proposed framework consists of an effective feature-based knowledge distillation structure by leveraging additional prediction head and a novel data augmentation method based on class-activation-map to address both catastrophic forgetting and class-imbalance issues simultaneously. Our method shows improved performance on all datasets, Food101-LT, VFN-LT, VFN-INSULIN and VFN-T2D, comparing with existing work. Finally, we conduct an ablation study to evaluate the effectiveness of each individual component and discuss the trade-off of potential techniques to achieve higher accuracy including the memory buffer capacity and transfer learning. Our future work includes designing an exemplar-free long-tailed continual learning framework. If successful, this could address the current challenges associated with large memory buffer and potential privacy concerns associated with storing food images.

### v2

@@ VI Conclusion
In this work, we focus on visual food recognition in long-tailed continual learning. We create an expanded dataset VFN186 and its three benchmark long-tailed food image datasets that exhibit the real-life food consumption frequency. The proposed end-to-end framework combines effective feature-based knowledge distillation and a novel data augmentation module, capable of learning new food classes in long-tailed data distribution without forgetting the learned knowledge. Our method outperforms existing approaches on all mentioned datasets. Future work includes developing an exemplar-free framework to tackle issues related to large memory buffers and privacy concerns with stored food images.